1. Given a 7.0x to 8.0x 2008E EV/EBITDA multiple range and 2008E EBITDA of $250 million, calculate an implied enterprise value range.
Answer #1

implied enterprise value (lower side of the range)=250*7=1750 million

implied enterprise value (upper side of the range)=250*8=2000 million

so  an implied enterprise value range is 1750 million to 2000 million

the above is answer..